# A quantitative analysis of Galilei’s observations of Jupiter satellites from the Sidereus Nuncius

### v1

@@ 3.4 Association of the satellites using a sky simulator
Figure 6: Observation 7, 15 January 1610, 19.55 CET: Galilei (top) vs simulation (bottom).
Each sketch was put side by side with the output of the Stellarium web-based sky simulator after taking care to see the values to those that were translated in modern time. All the images from Stellarium are shown together with the sketches of Galileo in Appendix A.
The output of the associations is given in Tabs. 5 and 6. In several cases two satellites were too close to be separated by Galileo. When this happens we decided to perform the association with one of the two. In general we find an almost perfect correspondence between the notes of Galileo and the simulator.
@@ 64 CGEI $\star$
In total the satellites that were potentially visible but are not recorded are 33 (Io 17 times, Europa 8 times, Ganymede 4 times and Callisto 4 times). We have excluded the cases in which the satellites were transiting the disk of Jupiter or were behind it. The reason for the inefficiency of the recordings is almost always due to the proximity of the satellite to the disk of the planet that must have been blurry. We can estimate that Galileo could see the satellites is they were away from the limb of the planet by at least approximately 2.5 diameters or about 1’40”. It should be noted that on January 7 1610 Jupiter was 4.28 AU away corresponding to about 45” for the disk angular diameter while on March 1 it was 5.02 AU away corresponding to 38”. This change in angular diameter of the system by 18% is not taken into account in the following analysis as we always report the elongations of satellites normalised to the angular diameter of the disk. Indeed the result suppor the validity of this assumption.
Observation 46 is probably one of the only that is difficult to reconcile with the prediction of the simulator. Galilei sees two satellites on the left and two on the right. It seems that either Io or Europa are on opposite sides of the disk while the simulator predicts both being on the right. We have checked if this could be related to a mismatch in the recorded hour but in the night of Feb 11 the pattern was the same as observation 45 (both Io and Europa left). The transit of both happens almost parallelly during the daytime of Feb 12 and during the early night of 12 they are both on the right side.
@@ 3.5 Periodograms of the satellite-tagged datasets
In this case the peaks of each satellite emerge without ambiguity at the expected position. Only for Io there is a less-significant, but still important, peak at about 1000 rad/y that had already been observed in the inclusive analysis of Fig. 5. In general the smaller peaks do not occur in correspondence of the expectations for other satellites supporting the fact that the rate of wrong-association is likely very low.
@@ 3.6 Sinusoidal fits and orbital parameters extraction
The elongation datasets, ($x_{i}$, $t_{i}$), separated by satellite have been fitted with a function of the form $x(t)=A\sin(\omega(t-t_{0})+\phi_{0})$ using the ROOT [6] fitter libraries. We have set the uncertainty on the point to match the r.m.s. of the residuals in practice setting the to the errors determined by the fit itself. The data with the superimposed best fita are shown in Fig. 8.
Figure 8: Fitted data points shown separately for each dataset associated to individual satellites. From top to bottom Io, Europa, Ganymede and Callisto. The units on the $x$-axis are fraction of years starting from 1/1/1610 and in the $y$-axis the elongation of the satellites in GJD units.
The same results are shown all together in Fig. 9.
Figure 9: Fitted data points shown separately for each dataset associated to individual satellites: Io (red), Europa (gray), Ganymede (orange) and Callisto (magenta). The units on the $x$-axis are fraction of years starting from 1/1/1610 and in the $y$-axis the elongation of the satellites in GJD units.
@@
The residuals of the fit ($x_{i}-x(t_{i})$) are shown for each satellite in Fig. 10. The top row shows them as a function of the point time while the bottom rows shows their distribution fitted with a Gaussian model.
@@
The model is reproducing the data very well without particular trends in the residuals or anomalies. In terms of the Jupiter disk in Galilei’s sketches (GJD) The r.m.s. of the residuals is 0.5 GJD for Io, 0.4 GJD for Europa, 0.56 GJD for Ganymede and 0.63 GJD for Callisto. Considering an average angular disk size of 42” during the measurements, and a factor 2.5 between the real and reported angular size, the previous r.m.s. correspond to values ranging from 40” to about 70”. This can be considered as an estimate of the pointing accuracy of the observations.
In the following we try to imagine how these data can be used to test the validity of Kepler third law for the Jupiter system and the 1:2:4 Laplace resonance between the periods of Io, Europa and Ganymede.
To do this test we have converted the radiuses of satellites’ orbits with a scaling factor chosen such that the scaled value for Ganymede corresponds to the modern measured value of 1.0704 million of km [^7] The scaling factor is 1.0704 Gm/$(4.482\pm 0.084)$ GJD = 2.39$\times$108 m/GJD..
@@
Figure 11 and Tab. 7 show a comparison between the values of the satellites’ orbits radiuses $A$ and their periods $T=2\pi/\omega$ as found by Galilei (bullets) and by modern determinations (histogram).
quantity Galilei rel. err. (%) stat.
$T$ (Io) (1.7730$\pm$ 0.0018) days 0.10
$T$ (Europa) (3.5539 $\pm$ 0.0037) days 0.10
$T$ (Ganymede) (7.151$\pm$0.012) days 0.16
$T$ (Callisto) (16.678 $\pm$ 0.051) days 0.31
$A$ (Io) (0.443 $\pm$ 0.019) Gm 4.3
$A$ (Europa) (0.687 $\pm$ 0.012) Gm 1.7
$A$ (Ganymede) (1.070 $\pm$ 0.020) Gm 1.9
$A$ (Callisto) (1.6482 $\pm$ 0.020) Gm 1.9
Table 7: Comparison of $T$ and $A$ obtained with this analysis and the modern determinations. Distances are in units of million km and times in days.
@@
On the $x$-axis we show the satellite identifier (from left to right Io, Europa, Ganymede, Callisto). Periods are in seconds, are shown in the top left plot while radiuses (in meters) in the top right plot. The bottom plots show the ratio between the measured value by Galilei and the modern value. The periods are compatible with modern values within their errors that range from 0.1% to 0.3%. As far as the radiuses are concerned the ratios are compatible with unity with 1-2 $\sigma$ for Io and Europa. The one of Ganymede is one by construction while for Callisto $A$ is underestimated by about 12%. The statistical errors on these parameters range from 1.9% for the outer satellites to 4.3% for Io. There is hence a trend for a significant underestimation of the major semi-axes for Callisto (See Fig. 12, bottom right plot).
Figure 12 (left) shows a scatter plot of $A$ vs $T$. Superimposed a curve of the type $T^{\frac{2}{3}}$. The curve is the function: $A(T)=k_{J}T^{\frac{2}{3}}=\left(\frac{GM_{J}}{4\pi^{2}}\right)^{\frac{1}{3}}T^{\frac{2}{3}}$ with $G=6.67430\times 10^{-11}$ Nm2kg-2 and $M_{J}=1.898\times 10^{27}$ kg hence $A(T)=5.02\times 10^{5}\times T^{\frac{2}{3}}$.
Figure 12: Kepler third law for the Jupiter system.
@@
In Fig. 13 we compute the ratios $T_{Eu}/T_{Io}-2$ and $T_{Eu}/T_{Io}-4$ to verify the level of accuracy in determining the orbital resonance effects. Black filled bullets indicate the measurements of Galilei and the blue hollow ones the modern values. Good agreement is found. Deviations from an exact 1:2:4 ratios, also present in modern data, are due to the slow precession of Io’s orbit’s closest point to Jupiter (perijove).
Figure 13: Test of the 1:2:4 Laplace resonance for Io, Europa and Callisto.
3.7 The performances of the Sidereus Nuncius “cannocchiale”
Unfortunately the instrument used in early 1610 has been lost but we have other instruments that were used by Galilei. Some optical tests performed on two of these instruments is available at [8] and considerations on its aberrations in [9]. In this section we want to imagine what could have been the performances of the telescope based on the accuracy obtained on the Jupiter system. In the following we use sharp images and apply a very simple Gaussian smearing plus a deformation that brings objects close to the edge of the field of view to appear closer to the center than they really are. This is inspired by the observation that orbits of Ganymede and Callisto appear to be underestimated. This should be verified with an optical simulation or with a system that tries to reproduce the original devices. Furthermore the effect could be also simply related to the increased difficulty of correctly estimating the angular separation when the two points have a larger separation separated. Also it is likely that the distances were estimated in steps, and that Jupiter might have not always been in the center of the field. Despite of this we have nevertheless included this effect in the simulation. The idea is to try and get an impression of what would have been the image available to Galilei.
In Fig. 14 we show how the two effects act on the Jupiter system.
Figure 14: Top view is the ideal image of the Jupiter system. Center picture introduces spherical aberration while the bottom one also adds in the effect of resolution.
Applying the same smearing and aberration effects to terrestrial targets allows to get an insight of the impression that the observation with the “cannocchiale” might have given to people of those times. We know that in August 1609, i.e. about five months before the observations of Jupiter, Galilei had demonstrated the potentials of the cannocchiale to the Doge of Venice, Leonardo Donà (Fig. 15, left), probably with a telescope that was still not as good as the later one.
Here we assume that the observations were done from the Campanile of San Marco and, for simplicity, we use the performances of the telescope of January 1610. As a target we have chosen a quite peculiar on: the Scala Contarini del Bovolo that is a very distinctive building located at 360 m East of the observation point and with a diameter of about 6 m (Fig. 15, center and right). It would have surely attracted the attention of the viewers from the Campanile.
Figure 15: Left: Galileo and the Doge (Bertini’s fresco). Right: Scala Contarini del bovolo from the Campanile of San Marco. Campanile of San Marco.
The close up on the right of Fig. 15 is given to show the, very narrow, field of view of the Galileian telescope (circle) that we estimate in about 20” or a third of degree.
The view of the scala is shown in a modern detailed picture in Fig. 16 (left) while the other pictures show the effect of resolution (center) and resolution and aberration (right).
Figure 16: Simulation of the possible view of the Cannocchiale when watching scala Contarini del Bovolo from the Campanile of San Marco. The picture on the left is at high resolution. The middle one reproduces the limited resolution while the right one, both resolution and spherical aberration. The same smearing and aberration parameters used in Fig. 14 have been applied.
Despite the distortions and the narrow field of view it is clear how amazing this could have been and indeed we know how enthusiastically this technological advance was received, granting Galilei an immediate increase in salary.

### v2

@@ 4 Association of the satellites using a sky simulator
Figure 12: Observation 7, 15 Jan. 1610, 19.55 CET: drawing vs simulation.
Each sketch was put side by side with the output of Stellarium after taking care to see the values to those that were translated in modern time. All the images from Stellarium are shown together with the sketches of Galileo in Appendix A. The observations 5.1, 40.1, 40.2, 42.1, 46.1, 52.1, 53.1, 61.1, 64.1 i.e. those where we only have the angular recordings but not the sketches are compared to the simulator in Appendix  B.
The output of the associations is given in Tabs. 6 and 7. In several cases two satellites were too close to be separated by Galileo. When this happens we decided to perform the association with one of the two. In general we find an almost perfect correspondence between the notes of Galileo and the simulator with the exception of a glitch for observation 46 of Feb 12 that will be discussed later.
@@ 4.1 Periodograms of the satellite-tagged datasets
In this case the peaks of each satellite emerge without ambiguity at the expected position. Only for Io there is a less-significant, but still important, peak at about 1000 rad/y that had already been observed in the inclusive analysis of Fig. 10. In general the smaller peaks do not occur in correspondence of the expectations for other satellites supporting the fact that the rate of wrong-associations is very low.
@@ 4.2 Sinusoidal fits and orbital parameters extraction
The elongations of dataset-1, ($x_{i}$, $t_{i}$), separated by satellite have been fitted with a function of the form $x(t)=A\sin(\omega(t-t_{0})+\phi_{0})$ using the ROOT [27] fitter libraries. For dataset-2 the model also accounts for a variation of the angular size of the Jovian system due to the farthening from Earth: $\theta(t)=\Theta(t)\sin(\omega(t-t_{0})+\phi_{0})$ where $\Theta(t)$ is a know function that can be obtained from the ephemerides. Indeed in the considered period (7 January-2 March) the distance of Jupiter passed from 4.28 to 5.03 AU and the phase passed from 100% to 99%. The angular diameter hence passed from 45.0 to 38.3 arcsec, i.e. Jupiter disk was 15% smaller at the end of the observation period. This effect does need to be taken into account for the analysis where the measurements are normalized to the Jupiter disk while it is relevant for the analysis in absolute angular distances as reported by Galilei. The variation of the equatorial diameter of Jupiter (from [1]) was fitted by a function of the form $\Delta(^{\prime})=46.2842-0.100186\times t-0.00103484\times t^{2}+9.86778\times 10^{-6}\times t^{3}$ with $t$ expressing the number of days since the first observation. We will later check if this model describes better the dataset-2 with respect to the simplified model.
We have tuned the uncertainty on the points to match the r.m.s. of the residuals in practice letting the errors determined by the fit itself by imposing to get a normalized $\chi^{2}$ of about one. The errors hence result as follows:
$\sigma_{I}$ $\sigma_{E}$ $\sigma_{G}$ $\sigma_{C}$
dataset-1 0.49 0.66 0.59 0.71
dataset-2 0.99 0.86 1.02 1.44
Table 8: Errors associated to the measurements in the sinusoidal fits.
This can be considered as an estimate of the pointing accuracy of the observations.
The data with the superimposed best fit are shown in Fig. 14 (dataset-1) and 15 (dataset-2).
Figure 14: Fitted data points from dataset-1 shown separately for each dataset associated to individual satellites. From top to bottom Io, Europa, Ganymede and Callisto. The units on the $x$-axis are fraction of years starting from 1/1/1610 and in the $y$-axis the elongation of the satellites in GJD units.
Figure 15: Fitted data points from dataset-2 shown separately for each dataset associated to individual satellites. From top to bottom Io, Europa, Ganymede and Callisto. The units on the $x$-axis are fraction of years starting from 1/1/1610 and in the $y$-axis the elongation of the satellites in GJD units.
The same results are shown superimposing all satellites together in Fig. 16 for dataset-1 and dataset-2.
Figure 16: Fitted data points from dataset-1 (top) and dataset-2 (bottom) shown separately for each dataset associated to individual satellites: Io (red), Europa (gray), Ganymede (orange) and Callisto (magenta). The units on the $x$-axis are fraction of years starting from 1/1/1610 and in the $y$-axis the elongation of the satellites in GJD units.
@@
The residuals of the fits ($x_{i}-x(t_{i})$, $\theta_{i}^{corr}-\theta(t_{i})$) are shown for each satellite in Fig. 17 as a function of time for the two datasets.
@@
The distribution of the residuals fitted with a Gaussian model is shown in Fig. 18. In this case we have multiplied the dataset-1 by 1.825 (see Fig. 6) in order to make them more easily comparable.
Figure 18: Distributions of the residuals for sinusoidal fits for Io (red), Europa (gray), Ganymede (orange) and Callisto (magenta) for the two datasets after scaling GJD units to primes.
These distributions show that for Io the dataset-1 seems significantly better fitted than dataset-2. It should be noticed anyway that the gaussian fit is not particularly good at modeling the data so this effect might appear overestimated. As we move to larger elongations the difference is moderate but the dataset-1 tends to be fitted a bit better. The reason for this is not easy to imagine and it would be important to know the relation between the notes of Galileo and the sketches in the Sidereus. We can suppose that in the printing process the drawings were drawn starting from the angular measurements (except for the first measurements). Still it is interesting to notice that dataset-1 seems significantly less discretized than the dataset-2 (see Fig. 6).
The models are reproducing the data very well without particular trends in the residuals or anomalies. The sigma of the gaussian fits are (0.3, 1.0, 1.0, 1.2) for dataset-1 and (0.8, 0.9, 1.5 and 1.8) for dataset-2. The fitted values with uncertainties are given in Tab. 9.
As we said we converted the observation times to modern hours in the analysis whenever available. It is interesting to cross check if this correction is effective. As a test we have tried to assume that all the measurements were done at 19.00 CET. Indeed as expected, we observe a very significant increase in the $\chi^{2}$ in the latter case, especially for Io for which this correction is generally quite important. Applying the correct observation the $\chi^{2}$ changes as follows: from 180.9 to 106.5 for Io, from 181.6 to 164.7 for Europa, from 155.0 to 158.4 for Ganymede and from 195.6 to 184.7 for Callisto.
4.3 Compatibility of datasets
As we have seen in Fig. 8 the two datasets are very close. Here we evaluate how compatible are the periods and phases of the sinusoids extracted from the fits. The results are shown in Fig. 19. As expected the extracted values are compatible to a very good level with the exception of Europa where there is a mild tension between the two datasets.
Figure 19: Confidence regions for $\omega$ and $\phi$ for dataset-1 and 2.
4.4 Effect of varying distance
The result of the fits with variable or fixed amplitude for Callisto dataset-2 is shown in Fig. 20. A better description is achieved with the fixed amplitude fit. The $\chi^{2}$ is consistent in the two fits but slightly lower for the fixed amplitude (about 4 units of $\chi^{2}$). This supports the fact that also the absolute measurement are in reality relative to the size of Jupiter.
Figure 20: Difference in the fit of the Callisto points for dataset-2 with the fixed amplitude and variable amplitude fit.
4.5 Comparison with ephemerides for the angular dataset
So far the elongations have been treated as a free parameter but in the case of dataset-2 we can check if the fitted elongations agree with the absolute prediction provided by the ephemerides[1]. The comparison of the data with the ephemerides is shown in Fig. 21.
Figure 21: Comparison between dataset-2 and the ephemeris of Jupiter providing the absolute elongations.
The data appear to be in excess by a significant fraction. To evaluate this factor we have compared these absolute predictions with the fit function with variable width. These two curves are compared in Fig. 22.
Figure 22: Comparison of the absolute predictions (dashed line) from the ephemerides and the fit (solid line) for dataset-2.
The fit and ephemerides are perfectly matching in terms of periods and initial phase. To match the amplitudes we had to scale the ephemerides down by a factor of (0.68, 0.65, 0.69, 0.73) for the four satellites and the result is shown in Fig. 41 in appendix. This result matches with the findings of Roche [6] that quotes a factor 1.4 (quoting the the inverse).
According to Giovanni Alfonso Borelli the elongations were estimated using a “micrometer” i.e. a scale that could slide along the tube of the telescope until the apparent size of a division, seen with the other naked eye, matched the diameter of Jupiter. In this case the overestimation might be due to the assumption that the angular size of Jupiter was 1 wide, hence 70% larger than the real value of 38-45′′. In this case these measurements would also be relative to the diameter of Jupiter and then it is not surprising that the amplitude of the orbits seems to be constant in the data. On the other hand in the Sidereus Galilei describes how the absolute measurement could be absolutely calibrated with terrestrial objects of known distance and dimensions. Probably the 30% discrepancy might due to a systematic uncertainty in the calibration or simply a deliberate choice dictated by the limited avaliable time that led him to choosing for Jupiter a diameter of one prime. It is interesting to notice that the estimation of the uncertainty diven by Galilei in the same section of the Sidereus is just one minute, or two. The estimation is in perfect agreement with our analysis (see f.e. Tab. 8).
Roche proposes a different explanation related to the fact that the measurement might have used the knowledge of the field of view of the telescope and to the fact that this field of view depends ([29]) on the aperture of the eye pupil. The pupil adapts to smaller diameters during daytime or when trying to make a calibration on the Moon hence producing a different f.o.v. with respect to the one that is achieved when observing Jupiter. This explanation is indeed intriguing and not completely easy to verify. Taking pictures naturally does not help as the effect of the pupil is not taken into account. It is certainly an interesting and clever interpretation that would deserve more scrutiny. On the other hand the fact that the fitted amplitudes do not show hints for reducing during towards the end of the considered period supports that the measurement are indeed relative to the perceived angular size of Jupiter. This explanation is discarded in by Roche ([6], pag. 25) on the basis that according to him Galilei estimated 2 for the size of the disk of Jupiter. This would make the overestimation too gross and fail this interpretation. Our analysis (Fig. 6 and Fig. 7) indeed indicates that Galileo was assuming a value of about 1 which is indeed about a factor 1.4 larger than the real angular extension.
4.6 Kepler III law and 1:2:4 resonance
[…]
@@ 4.6 Kepler III law and 1:2:4 resonance
Figure 23 shows a comparison between the values of the satellites’ orbits radiuses $A$ and their periods $T=2\pi/\omega$ as found by Galilei (bullets) and by modern determinations (histogram).
Satellite Value Uncertainty e.r. (%)
Dataset-1
AA (/J. drawn diameter) Io 1.830 0.064 3.5
AA (/J. drawn diameter) Europa 2.884 0.065 2.2
AA (/J. drawn diameter) Ganymede 4.342 0.064 1.5
AA (/J. drawn diameter) Callisto 6.985 0.069 0.98
$T$ [days] Io 1.7727 0.0015 0.087
$T$ [days] Europa 3.5551 0.0049 0.14
$T$ [days] Ganymede 7.1175 0.0094 0.13
$T$ [days] Callisto 16.6693 0.0395 0.24
Dataset-2
AA () Io 3.13 0.12 3.9
AA () Europa 5.242 0.086 1.6
AA () Ganymede 7.86 0.11 1.4
AA () Callisto 12.95 0.13 1.0
$T$ [days] Io 1.774 0.0018 0.10
$T$ [days] Europa 3.5415 0.0035 0.10
$T$ [days] Ganymede 7.1403 0.0095 0.13
$T$ [days] Callisto 16.613 0.043 0.26
Table 9: Fitted amplitudes (AA) and orbital periods ($T$) with uncertainties and relative errors for the four Galilean satellites.
@@
On the $x$-axis we show the satellite identifier (from left to right Io, Europa, Ganymede, Callisto). Periods are in seconds, are shown in the top left plot while radiuses (in meters) in the top right plot. The bottom plots show the ratio between the measured value by Galilei and the modern value. The periods are compatible with modern values within their errors that range from 0.1% to 0.3%. As far as the radiuses are concerned the ratios are compatible with unity with 1-2 $\sigma$ for Io and Europa. The one of Ganymede is one by construction while for Callisto $A$ is underestimated by about 10% for dataset-1 and 5% for dataset-2. The statistical errors on these parameters range from 1.9% for the outer satellites to 4.3% for Io. There is hence a trend for a significant underestimation of the major semi-axes for Callisto (See Fig. 24, bottom right plot).
Figure 24 (left) shows a scatter plot of $A$ vs $T$. Superimposed a curve of the type $T^{\frac{2}{3}}$ [^5] The curve is the function: $A(T)=k_{J}T^{\frac{2}{3}}=\left(\frac{GM_{J}}{4\pi^{2}}\right)^{\frac{1}{3}}T^{\frac{2}{3}}$ with $G=6.67430\times 10^{-11}$ Nm2kg-2 and $M_{J}=1.898\times 10^{27}$ kg hence $A(T)=5.02\times 10^{5}\times T^{\frac{2}{3}}$..
Figure 24: Left: Kepler third law for the Jupiter system. Right: Test of the 1:2:4 Laplace resonance for Io, Europa and Callisto.
@@
In Fig. 24 (right) we compute the ratios $T_{E}/T_{I}-2$ and $T_{E}/T_{I}-4$ to verify the level of accuracy in determining the orbital resonance effects. Black filled bullets indicate the measurements of Galilei and the blue hollow ones the modern values. Good agreement is found. Deviations from an exact 1:2:4 ratios, also present in modern data, are due to the slow precession of Io’s orbit’s closest point to Jupiter (perijove).
5 Efficiency near the disk and separating power
In Fig. 25 we show the distribution of the positions of satellites that went undetected probably due to being too close to the disk of Jupiter. There is a sort of indication of a lower efficiency for the western side. This is the side that sits in the field of view of the telescope for less time due to Earth’s rotation. It might still be a statistical fluctuation.
Figure 25: Map of satellites that went undetected probably due to being too close to the disk of Jupiter. The disks in gray represent that are less reliable (obs. 23, 38, 39, 48) as the hour is not specified hence the satellite might have been closer to Jupiter than what is assumed in the simulator (observation at one hour after sunset).
In Fig. 25 we show the distribution of the positions of pairs of satellites that were not resolved into a single object. The largest unresolved case has an angular separation of 31-37′′ but other typical cases are in the order of 15-17′′. This estimate is somewhat affected by the uncertainty in the exact time of the observation, especially when Io was involved or in the proximity of Jupiter where the velocity is higher. It is interesting to notice that the pairs in cyan color and dark blue, were observed as a single object even if they fall in the region where efficiency was low (see Fig. 25).
Figure 26: Map of pairs of satellites that were perceived as a single object.
There are then the cases where the satellites were quite close to the diameter but were anyway detected. This happened up to angular distance distances of about one equatorial diameter or less from the limb (see observation 5 and 18, both with a double pair unresolved, and 32, 45 with a single satellite quite close but seen).
For observation 32 Galileo mentions a distance from the limb of 20 arcsec which corresponds to about half disk. The dimension of the disk reported in the drawing in that case is similar to the real one. This makes us thing that the problem for the loss of efficiency was likely more related to the glare of the planet rather than an enlarged disk due to aberrations. It should be noted that in a publication of 1992 Greco et al. [30] showed that Galileo’s lenses, though simple (a positive objective and a negative eyepiece), were made with remarkable skill for the early 17th century. The objective lenses showed near-diffraction-limited quality, with errors only about 1/4 wavelength. Eyepieces were of poorer quality, but their small aperture angles minimized the effect on image quality. Aberrations were limited by applying collimators to the objective also using elliptical profiles (…that the opening left uncovered should be oval in shape, for in this way the objects will be seen much more distinctly.”) [18, 31] (pag. 278).
In summary the efficiency is practically zero when the satellite is less than one diameter from the limb, it rises then above 1 diameter but there is still a significant inefficiency until more than two diameters from the limb, especially on the west side.
We have finally performed a more quantitative evaluation of the dependency of efficiency on the distance from Jupiter. We have taken the sinusoidal fits of the four satellites and evaluated the elongations at the time of each observation for all four satellites. The distribution of the absolute value of the elongations is shown in the black histogram of Fig. 27. The red histogram shows the same predictions only taken when the satellite was actually observed according to the pattern described in Tab. 6 and 7. The efficiency is obtained by taking the ratio of the two distributions and is shown in the top right plot of the same Figure as a function of the predicted elongation of the satellite. The drop of efficiency due to the proximity to the disk is very evident and shows a progressive trend as the distance is lower. The inefficiencies at higher distances are basically due to the fact that sometimes satellites are too close so the efficiency goes to 50% when they are not resolved.
Figure 27: Top left: distribution of all the predicted elongations (black histogram) and of those that correspond to an observation for dataset-1. Top right: detection efficiency obtained as the ration between the red and black histogram. Bottom, same as top but for dataset-2.
There are cases where two very close satellites were resolved. For example Io and Europa are at about one Jupiter diameter from each other for observation 37 or in observation 48 they are at about 3/4 of the equatorial diameter or Ganymede and Europa that in observation 64 are again at about 3/4 of the diameter (i.e. $\sim$. 30′′ being this observation done a the beginning of March when the diameter was about 30 arcsec).
In total the satellites that were potentially visible but are not recorded are 33 (Io 17 times, Europa 8 times, Ganymede 4 times and Callisto 4 times). We have excluded the cases in which the satellites were transiting the disk of Jupiter or were behind it. The reason for the inefficiency of the recordings is almost always due to the proximity of the satellite to the disk of the planet that must have been blurry. We can estimate that Galileo could see the satellites is they were away from the limb of the planet by at least approximately 2.5 diameters or about 1’40”. It should be noted that on January 7 1610 Jupiter was 4.28 AU away corresponding to about 45” for the disk angular diameter while on March 1 it was 5.02 AU away corresponding to 38”. This change in angular diameter of the system by 18% is not taken into account in the following analysis as we always report the elongations of satellites normalized to the angular diameter of the disk. Indeed the result support the validity of this assumption.
Observation 46 is probably one of the only that is difficult to reconcile with the prediction of the simulator. Galilei observes right after sunset and sees two satellites on the left and two on the right. It seems that either Io or Europa are on opposite sides of the disk while the simulator predicts both being on the right. We have checked if this could be related to a mismatch in the recorded hour but in the night of Feb 11 the pattern was the same as observation 45 (both Io and Europa left). Even assuming an observation in the morning of 12/2 does not explain the anomaly. The transit of both happens during the daytime of Feb 12 and during the early night of 12 they are both on the right side. It is interesting to notice that this anomaly, independently observed in this work, it had been already noticed in the literature (Drake, Levi/Levi-Donati, Bernieri). Galileo was traveling back from Venice to Padua that day, the implication being that he might have been hasty due to the constraints of the travel, or maybe drew the sketch some time later, from memory. This explanation seems very likely. Indeed we think that a close-by moon was there (Europa) and it was just recorded at about the correct elongation but on the wrong side. What remains strange is that 3h 20’ later (at the fifth hour) the satellite on the East had disappeared while in principle Europa had become more visible to the West.